# Refinable modeling for unbinned SMEFT analyses

### v1

@@ 4.4.3 SMEFT modeling
The SMEFT effects can be modeled with synthetic data, initially motivating the development of optimal SMEFT observables [13, 14, 15, 16, 17, 18]. The ME in Eq. 14 can efficiently be recomputed as a function of the POIs ${\boldsymbol{\theta}}$. The SMEFT dependence of the DCR is
@@
where we have dropped the nuisance parameter dependence because the numerator and the denominator are evaluated for the same ${{\boldsymbol{\nu}}_{0}}$. The conditional probabilities in the third term, which are not tractable, cancel in the ratio. The remainder is the parton-level DCR, which is available at the level of the ME generator; the ratio $r(\boldsymbol{x}_{i},\boldsymbol{z}_{i}|{{\boldsymbol{\theta}}},{\boldsymbol{\theta}_{0}})$ does not depend on $\boldsymbol{x}_{i}$. By computing the per-event polynomial coefficients $r^{(m)}(\boldsymbol{z}_{\textrm{p},i})$ and $r^{(mn)}(\boldsymbol{z}_{\textrm{p},i})$ with $m,n=1,\ldots,N_{{{\boldsymbol{\theta}}}}$ from the ME-squared terms evaluated for a sufficient number of different ${\boldsymbol{\theta}}$, we can obtain a parametrization of $r(\boldsymbol{x}_{i},\boldsymbol{z}_{i}|{{\boldsymbol{\theta}}},{\boldsymbol{\theta}_{0}})$ valid in the whole SMEFT parameter space. Synthetic data sets $\mathcal{D}({{\boldsymbol{\theta}}},{{\boldsymbol{\nu}}_{0}})$ are, therefore, available. We provide the technicalities in Appendix A. We also note that the SM point in the denominator hypothesis is not exceptional. Indeed, we could write Eq. 36 for general ${\boldsymbol{\theta}_{0}}$, offering the possibility to simulate at EFT parameter points other than the SM.
@@ 4.5 Large sample limit and overflow bins
Learning surrogates of the likelihood ratio requires sufficient simulated data, and we shall assume that the large sample limit is respected when minimizing loss functions in the following sections. For any finite simulated data set, the observables with energy units typically imply a threshold beyond which the simulation is too scarce. Because SMEFT effects can grow with energy, removing events with features in the tails of the distribution of energetic variables is generally detrimental. Instead, for each such variable in $\boldsymbol{x}$, we can define a threshold beyond which we do not trust the modeling of the differential cross-section but still have acceptable uncertainties in the cumulative yield. We thus accumulate all events with features above certain thresholds in several “overflow bins” (OF) and treat these bins with standard Poisson likelihoods. Concretely, suppose $x_{n}$ is a feature in the vector of observables $\boldsymbol{x}^{\prime}$ representing the event before the overflow is introduced. In that case, we need to deal with the fact that synthetic data insufficiently samples $p(\boldsymbol{x}^{\prime}|x_{n}>x_{n,0},{{\boldsymbol{\theta}}},{\boldsymbol{\nu}})$. The simplest solution is to drop all event features for $x_{n}>x_{n,0}$ and represent the event solely by the fact that it is in the overflow of $x_{n}$, i.e.,
@@
For a number of overflow bins ($N_{\textrm{OF}}$), one for each observable where it is needed, we thus reduce the observation to counting variable $N(\textrm{OF}_{n})$ instead of the fully unbinned information. The number $N_{\textrm{OF}}$ characterizes the measurement, while the observed number of events in the n-th overflow bin $N(\textrm{OF}_{n})$ characterizes the observation. Explicitly, $\mathcal{D}^{\prime}=\mathcal{D}\cup\{N(\textrm{OF}_{n})\}_{n=1}^{N_{\textrm{OF}}}$. The likelihood associated with the observation in the overflow bins follows from Eq. 5 as a Poisson factor, one for each overflow bin, as
@@
As per the assumption, the predicted yield in the overflow bin as a function of the POIs and the nuisances, denoted by $\lambda(\textrm{OF}_{n}|{{\boldsymbol{\theta}}},{\boldsymbol{\nu}})$, is available from simulation with sufficient precision. The extra factor is a standard binned likelihood term that should be treated no differently than in standard binned SMEFT data analyses. To simplify the formulae of the following discussions, we assume, from now on, that this factor is accounted for in Eq. 5 and, in turn, that $\boldsymbol{x}\in\mathcal{X}$ implies the event is not in any of the overflow bins.
@@ 5 Learning from simulation
With procedures in place for obtaining simulated and synthetic data sets, we next describe, in general, how we can learn parameterizations from commonly used loss functions, such as the cross-entropy loss. We work with a general expressive function $\hat{f}(\boldsymbol{x})$ but, for now, do not specify its implementation. It can be a neural network, a boosted tree algorithm, or any other trainable multivariate predictive function.
@@ 5.1 Likelihood-ratio trick and cross-entropy loss
We start with the well-known “likelihood-ratio trick”, the basis for the learning tasks in this work: A sufficiently expressive machine trained on a classification task learns (a monotonous function of) the likelihood ratio. If the training data is normalized according to the differential cross-section, the classifier learns the DCR.
This fact is at the heart of learning techniques for parametric surrogates. Let us take two fixed hypotheses (${\boldsymbol{\theta}_{0}}$, ${{\boldsymbol{\nu}}_{0}}$) and (${\boldsymbol{\theta}_{1}}$, ${{\boldsymbol{\nu}}_{1}}$), obtained either by independent simulation or produced synthetically as described in Sec. 4.4, and minimize the cross-entropy loss function
@@
Because $\hat{f}(\boldsymbol{x})$ does not depend on $\boldsymbol{z}$ by construction, we can formally integrate over the latent configuration using $\int{\textrm{d}}\boldsymbol{z}\,p(\boldsymbol{x},\boldsymbol{z}|{{\boldsymbol{\theta}}},{\boldsymbol{\nu}})=p(\boldsymbol{x}|{{\boldsymbol{\theta}}},{\boldsymbol{\nu}})$, separately for the summands in Eq. 39. The solution is analogously expressed in terms of latent-space integrals and reads
@@
This expression is a monotonous function of the desired DCR for two fixed choices of the model parameters. It can be rearranged to
@@
The simulation-based approximation of Eq. 39, suitable for a concrete implementation in computer code, is
@@
It is a different approximation for the same minimization problem. The main difference to Eq. 43 is the absence of independent stochastic fluctuations in the two terms. In both cases, $\hat{f}(\boldsymbol{x})$ is implemented as a finitely but sufficiently expressive ML algorithm, approximating the exact solution. We denote this approximation by
@@ 5.2 Machine-learning systematic parametrizations
For learning parameterizations, the idea is to substitute $\hat{f}$ with a suitable parametric ansatz that implements the ${\boldsymbol{\nu}}$ dependence. The loss function is summed over a set of model parameter “base points”, denoted by $\mathcal{V}$. The parametrization can be determined with synthetic data from enough base points. We drop ${\boldsymbol{\theta}}$ from the formulae because we will factorize this dependence in Sec. 6. We also chose the fully calibrated SM parameter point as the reference, ${{\boldsymbol{\nu}}_{0}}={\mathbf{0}}$. The ansatz
@@
eliminates the monotonous dependence from Eq. 41. The ML estimate of the DCR is then $\hat{S}(\boldsymbol{x}|{\boldsymbol{\nu}})=\exp(\hat{T}(\boldsymbol{x}|{\boldsymbol{\nu}}))$. The exponential function removes the necessity of ensuring that $\hat{S}(\boldsymbol{x}|{\boldsymbol{\nu}})$ must be positive by construction. Inserting Eq. 47 into Eq. 39, leads to
@@
The ellipsis indicates that cubic or higher terms could be added if needed. The functional form is equivalent to the ansatz in Refs. [22, 17]. Notably, we include the possibility that some of the coefficient functions are chosen to be absent.
@@
Each term in the sum in Eq. 52 contains two expectations; we emphasize that the first expectation in every term corresponds to the SM. The coefficient functions $\hat{\Delta}_{A}(\boldsymbol{x})$ can be implemented as neural networks or any other trainable multi-variate regressor. Because, in most cases, the factorization of systematic effects is an accurate simplification, the number of coefficient functions that must be learned simultaneously is relatively small. For the example of a single nuisance, linear or quadratic accuracy in the Eq. 53 is achieved by one or two coefficient functions, respectively, and sufficient in most cases. Higher-order terms can be included as necessary.
@@ 5.3 Two-point alternatives
No tractable simulation exists for several systematic effects, such as binary choices in the generator setup. In this case, we only can obtain two alternate simulations. The sum over $\mathcal{V}$ in Eq. 52 becomes trivial, simplifying the learning task. A single nuisance parameter $\nu_{2P}$ shall interpolate between the nominal choice, associated with $\nu_{\textrm{2P}}=0$, and the alternative, corresponding to $\nu_{\textrm{2P}}=1$. Only the linear approximation can be learned without predictions for more nuisance parameter values. Because $\nu_{\textrm{2P}}=1$ is the only value in $\mathcal{V}$, we have
@@
and $\hat{\Delta}(\boldsymbol{x})$ is a singe-valued coefficient function. Minimizing the loss function, we find the estimate for the DCR
@@
$\hat{S}_{\textrm{2P}}(\boldsymbol{x}|\nu_{\textrm{2P}})=\exp\left(\nu_{\textrm{2P}}\hat{\Delta}(\boldsymbol{x})\right)\simeq\left(\frac{{\textrm{d}}\sigma_{1}(\boldsymbol{x})}{{\textrm{d}}\sigma_{0}(\boldsymbol{x})}\right)^{\nu_{\textrm{2P}}}$ (55)
amount to, as expected, an $\boldsymbol{x}$-dependent linear interpolation of the logarithm of the DCR.

### v3

@@ 4.4.3 SMEFT modeling
SMEFT effects can be modeled with synthetic data, which initially motivated the development of optimal SMEFT observables [13, 14, 15, 16, 17, 18]. The ME in Eq. 14 can be efficiently recomputed as a function of the POIs ${\boldsymbol{\theta}}$. The SMEFT dependence of the DCR is
@@
where we have omitted the nuisance parameter dependence, as the numerator and denominator are evaluated for the same ${{\boldsymbol{\nu}}_{0}}$. The conditional probabilities in the third term, which are not tractable, cancel in the ratio with excellent accuracy. The remainder is the parton-level DCR, which is available at the level of the ME generator. Thus, $r(\boldsymbol{x}_{i},\boldsymbol{z}_{i}|{{\boldsymbol{\theta}}},{\boldsymbol{\theta}_{0}})$ does not depend on $\boldsymbol{x}_{i}$. By calculating the per-event polynomial coefficients $r^{(m)}(\boldsymbol{z}_{\textrm{p},i})$ and $r^{(mn)}(\boldsymbol{z}_{\textrm{p},i})$ with $m,n=1,\ldots,N_{{{\boldsymbol{\theta}}}}$ from the ME-squared terms at various ${\boldsymbol{\theta}}$ values, we can construct a parametrization of $r(\boldsymbol{x}_{i},\boldsymbol{z}_{i}|{{\boldsymbol{\theta}}},{\boldsymbol{\theta}_{0}})$ valid across the entire SMEFT parameter space, making synthetic data sets $\mathcal{D}({{\boldsymbol{\theta}}},{{\boldsymbol{\nu}}_{0}})$ readily available. We provide further details in Appendix A. Note that the SM point in the denominator is not unique; Eq. 36 can be applied for any ${\boldsymbol{\theta}_{0}}$, allowing for simulation at EFT parameter points other than the SM.
@@ 4.5 Large sample limit and overflow bins
Learning surrogates of the likelihood ratio requires sufficient simulated data, and in the following sections, we assume the large sample limit when minimizing loss functions. For any finite simulated data set, observables with energy units often imply a threshold beyond which simulation becomes too sparse. Since SMEFT effects can increase with energy, removing events in the tails of energetic variable distributions is generally counterproductive. Instead, for each such variable in $\boldsymbol{x}$, we can define a threshold beyond which we do not fully trust the modeling of the differential cross-section but still find acceptable uncertainties in the cumulative yield.
We accumulate all events with features above certain thresholds in several “overflow bins” (OF) and treat these bins using standard Poisson likelihoods. Suppose $x_{n}$ is a feature in the vector of observables $\boldsymbol{x}^{\prime}$ representing the event before introducing overflow. In that case, we must address the fact that synthetic data insufficiently samples $p(\boldsymbol{x}^{\prime}|x_{n}>x_{n,0},{{\boldsymbol{\theta}}},{\boldsymbol{\nu}})$. The simplest solution is to drop all event features for $x_{n}>x_{n,0}$ and represent the event only by its presence in the overflow of $x_{n}$, i.e.,
@@
For a number of overflow bins ($N_{\textrm{OF}}$), one for each required observable, we reduce the observation to the counting variable $N(\textrm{OF}_{n})$ instead of using fully unbinned information. The number $N_{\textrm{OF}}$ characterizes the measurement, while the observed number of events in the $n$-th overflow bin $N(\textrm{OF}_{n})$ characterizes the observation. Explicitly, $\mathcal{D}^{\prime}=\mathcal{D}\cup\{N(\textrm{OF}_{n})\}_{n=1}^{N_{\textrm{OF}}}$. The likelihood for the overflow bin observation follows from Eq. 5 as a Poisson factor for each overflow bin as
@@
As assumed, the predicted yield in each overflow bin as a function of the POIs and nuisances, denoted by $\lambda(\textrm{OF}_{n}|{{\boldsymbol{\theta}}},{\boldsymbol{\nu}})$, is available from simulation with sufficient precision. This additional factor is a standard binned likelihood term and should be handled as in traditional binned SMEFT analyses. To simplify the formulas in the following sections, we assume that this factor is included in Eq. 5 and that $\boldsymbol{x}\in\mathcal{X}$ implies the event is not in any overflow bin.
@@ 5 Learning from simulation
With procedures in place for obtaining simulated and synthetic data sets, we now outline how parametrizations can be learned using common loss functions, such as cross-entropy loss. We consider a general expressive function $\hat{f}(\boldsymbol{x})$, without specifying its implementation; it could be a neural network, the BPT from Sec. 5.4, or any other trainable multivariate predictive function.
@@ 5.1 Likelihood-ratio trick and cross-entropy loss
We begin with the well-known “likelihood-ratio trick,” which underpins the learning tasks in this work: a sufficiently expressive machine trained on a classification task learns a (monotonic function of) the likelihood ratio. If the training data is normalized by the differential cross-section, the classifier learns the DCR. This fact is at the heart of learning techniques for parametric surrogates. Let us take two fixed hypotheses (${\boldsymbol{\theta}_{0}}$, ${{\boldsymbol{\nu}}_{0}}$) and (${\boldsymbol{\theta}_{1}}$, ${{\boldsymbol{\nu}}_{1}}$), obtained either by independent simulation or produced synthetically as described in Sec. 4.4, and minimize the cross-entropy loss function
@@
Because $\hat{f}(\boldsymbol{x})$ does not depend on $\boldsymbol{z}$ by construction, we can formally integrate over the latent configuration using $\int{\textrm{d}}\boldsymbol{z}\,p(\boldsymbol{x},\boldsymbol{z}|{{\boldsymbol{\theta}}},{\boldsymbol{\nu}})=p(\boldsymbol{x}|{{\boldsymbol{\theta}}},{\boldsymbol{\nu}})$, separately for the summands in Eq. 39. The solution is then expressed in terms of latent-space integrals as
@@
This expression is a monotonous function of the DCR for two fixed choices of the model parameters and can be rearranged to
@@
Alternative loss functions and their minima are discussed in Appendix B. The simulation-based approximation of Eq. 39, suitable for a concrete implementation in computer code, is
@@
The main difference to Eq. 43 is the absence of independent stochastic fluctuations in the two terms. In both cases, $\hat{f}(\boldsymbol{x})$ is implemented as a finitely but sufficiently expressive ML algorithm, approximating the exact solution. We denote this approximation by
@@ 5.2 Machine-learning systematic parametrizations
To learn parametrizations, we replace $\hat{f}$ with a suitable parametric ansatz that captures the ${\boldsymbol{\nu}}$ dependence. The loss function is summed over a set of model parameter points (base points), denoted by $\mathcal{V}$. The parametrization can be determined using synthetic data from a sufficient number of base points. We omit ${\boldsymbol{\theta}}$ in the formulas, as we will factorize this dependence in Sec. 6. The fully calibrated SM parameter point serves as the reference, ${{\boldsymbol{\nu}}_{0}}={\mathbf{0}}$. The ansatz
@@
eliminates the monotonous dependence from Eq. 41. The ML estimate of the DCR is then $\hat{S}(\boldsymbol{x}|{\boldsymbol{\nu}})=\exp(\hat{T}(\boldsymbol{x}|{\boldsymbol{\nu}}))$. The exponential function removes the necessity of ensuring that $\hat{S}(\boldsymbol{x}|{\boldsymbol{\nu}})$ must be positive. Inserting Eq. 47 into Eq. 39, leads to
@@
The ellipsis indicates that cubic or higher terms can be added as needed, allowing to parametrize the systematic effects, in principle, with arbitrary precision. The functional form is equivalent to the ansatz in Refs. [22, 17]. Notably, we include the possibility that some of the coefficient functions are chosen to be absent.
@@
Each term in the sum in Eq. 52 contains two expectations; note that the first expectation in each term corresponds to the SM. The coefficient functions $\hat{\Delta}_{A}(\boldsymbol{x})$ can be implemented as neural networks or any other trainable regressor. In most cases, the factorization of systematic effects is a reliable simplification, so the number of coefficient functions that need to be learned simultaneously remains small. For a single nuisance, one or two coefficient functions are enough to achieve linear or quadratic accuracy in Eq. 53, which is usually sufficient. Higher-order terms can be added as needed.
@@ 5.3 Two-point alternatives
For some systematic effects, like binary generator choices, no tractable simulation exists–only two alternate simulations. In this case, the sum over $\mathcal{V}$ in Eq. 52 becomes trivial, simplifying the learning task. A single nuisance parameter $\nu_{2P}$ interpolates between the nominal choice ($\nu_{\textrm{2P}}=0$) and the alternative ($\nu_{\textrm{2P}}=1$). Without predictions for more values, we can only learn a linear approximation. With $\nu_{\textrm{2P}}=1$ as the sole value in $\mathcal{V}$, we get
@@
where $\hat{\Delta}(\boldsymbol{x})$ is a single-valued coefficient function. Minimization provides an estimate
@@
$\hat{S}_{\textrm{2P}}(\boldsymbol{x}|\nu_{\textrm{2P}})=\exp\left(\nu_{\textrm{2P}}\hat{\Delta}(\boldsymbol{x})\right)\simeq\left(\frac{{\textrm{d}}\sigma_{1}(\boldsymbol{x})}{{\textrm{d}}\sigma_{0}(\boldsymbol{x})}\right)^{\nu_{\textrm{2P}}},$ (55)
which is an $\boldsymbol{x}$-dependent linear interpolation of the logarithm of the DCR.